Please help I'm having trouble trying to solve logarithm equation log(y+2)=1-log (2y-6)
Answer by ikleyn(52905)

.
Please help I'm having trouble trying to solve logarithm equation log(y+2)=1-log (2y-6)
~~~~~~~~~~~~~~~~~~~~~

log(y+2) = 1 - log (2y-6)   --->  log(y+2) + log (2y-6) = 1  --->  (y+2)*(2y-6) = 10.


Can you solve this quadratic equation ?


When you complete solution, check if the roots suit under logarithms.